Two friends meet together at the bike path. One is on roller blades and the other bikes. The biker goes 8mph faster than the skater. The biker goes 52 miles and the skater goes 28 miles. How fast is each going?
Answer by jorel1380(3719)

52/x=28/x-8
52(x-8)=28x
52x-416=28x
24x=416
x=17.333
x-8=9.3333
The biker is going 17.333 mph. The skater is going 9.333 mph..